NvidiaIntely (nvidiaintely):

A rock has 12.5 percent of its original amount of potassium-40 remaining in it; potassium-40 has a half-life of 1.25 billion years. How long ago was the rock formed? 1.25 billion years ago 2.5 billion years ago 3.75 billion years ago 5 billion years ago

NvidiaIntely (nvidiaintely):

@Kevin

OpenStudy (kevin):

Nt = 12.5/100 No Nt/ No = 12.5/100 Nt/No = 1/2(t/ delta t) 12.5/100 = 1/2(t/1.25) t = 3.75

NvidiaIntely (nvidiaintely):

Thanks!